Write down energy in 2 dimensional for a particle.
1
Expert's answer
2020-11-02T09:25:42-0500

We can solve the Schrodinger problem for a particle in 2 - dimension hole:

"\widehat H \Psi_{n,m} = E_{n,m}\Psi_{n,m}"

where :

"\widehat H \rightarrow -\frac{\hbar^{2}}{2m}\left(\frac{\partial^2}{\partial x^2} + \frac{\partial^2}{\partial y^2}\right) + U(x,y)"

if we solve it for a hole with rectangle as a shape, then we get the level of energy fo such system:

"E_{n,m} = \frac{\pi^2\hbar^2}{2m}\left(\frac{n^2}{a^2} + \frac{m^2}{b^2}\right)"

n,m - quantum numbers

a,b - sides of the rectangle